# DP-Fusion: Token-Level Differentially Private Inference for Large Language Models

### v3

@@ Abstract
Large language models (LLMs) do not preserve privacy at inference-time. The LLM’s outputs can inadvertently reveal information about the model’s context, which presents a privacy challenge when the LLM is augmented via tools or databases containing sensitive information. Existing privacy-preserving methods at inference-time have significant limitations since they (i) lack provable guarantees or (ii) have a poor utility/privacy trade-off. We propose DP-Fusion, a Differentially Private Inference (DPI) mechanism for LLMs that provably bounds the influence a set of tokens in the context can have on the LLM’s output. DP-Fusion works as follows: (1) label a subset of sensitive tokens, (2) infer the LLM without any sensitive tokens to obtain a baseline, (3) infer the LLM with the sensitive tokens, and (4) blend distributions so that the final output remains within a bounded distance of the baseline distribution. While this per-token influence bound also mitigates jailbreak-style prompt injection, we focus on document privatization, where the goal is to paraphrase a document containing sensitive tokens, e.g., personally identifiable information, so that no attacker can reliably infer them from the paraphrased document while preserving high text quality. The privacy/utility trade-off is controlled by $\epsilon$, where $\epsilon=0$ hides sensitive tokens entirely, while higher values trade off privacy for improved text quality. We show that our method creates token-level provably privatized documents with substantially improved theoretical and empirical privacy, achieving $6\times$ lower perplexity than related DPI methods. Our code and data are publicly available here: https://github.com/MBZUAI-Trustworthy-ML/DP-Fusion-DPI.
@@ 1 Introduction
Figure 1: An overview of DP-Fusion for differentially private LLM inference whose output is revealed to a potentially untrusted user. Sensitive tokens are redacted PII.
@@
Large Language Models (LLMs) are trained once and deployed many times. During deployment, LLMs process unseen data they were not trained on, such as user prompts, tool calls or external databases. A privacy challenge emerges when data contains sensitive information such as passwords or Personally Identifiable Information (PII) (inference_time_1; regulation2016regulation) that the LLM must not reveal to a user. Consider a hospital that wants to deploy LLMs to assist users in matching their symptoms to historical records from a large document dataset. Many users contributed their doctor’s notes, including health details such as a disease history or a treatment plan linked to PII (pii_2), but expect privacy against re-identification.
@@
However, deploying an LLM introduces unique privacy risks since generated tokens could inadvertently and silently leak sensitive data. The challenge is protecting sensitive data while maintaining high service quality. A straightforward solution would be to carefully label sensitive tokens and scrub them from all documents. Scrubbing is widely applied in practice, but overly aggressive scrubbing has been shown to severely harm utility (pii_1). A better solution could be to fully re-write documents for privacy, e.g., through paraphrasing (mattern2022limits). However, doing inference naively (i) lacks provable guarantees and (ii) our experiments show that attackers can still reliably infer sensitive information when they know which model was used to create the paraphrased text. Private inference solutions fall into two categories: (a) modifying the LLM’s context (e.g., via scrubbing) or (b) modfying the inference process. Dataset-based techniques include randomized methods to replace sensitive tokens in the input (custext; rantext; santext). Inference-based techniques modify the model or inference process, e.g., via fine-tuning, prompt engineering (staab2024large), adding noise to the output distribution (dpdecod; utpala-etal-2023-locally). However, existing dataset and inference-based approaches achieve poor privacy/utility trade-offs, either by over-sanitizing the input or by providing weak or no formal guarantees.
@@
We introduce DP-Fusion, a token-level Differentially Private Inference (DPI) method for LLMs that provably bounds the influence of sensitive tokens in the context on generated tokens in its output. Figure˜1 illustrates an overview of our method, which computes the next token to a query on a document containing PII. We first remove the PII from the document, then run the LLM on both the original document (with PII) and the redacted version (without PII). Finally, we mix the probability distributions from both runs, so that the distance between the mixed and original distributions is bounded, sample the next token and return it to the user. Crucially, the attacker’s advantage at inferring the secret is provably bounded even if the query is chosen adversarially (e.g., by selecting a jailbreak attack (wei2023jailbroken)). We empirically demonstrate that our method substantially outperforms all surveyed DPI methods in the utility/privacy trade-off.
@@ 2 Background
An LLM has a context, which typically includes a (i) system prompt, (ii) user queries, (iii) LLM responses and (iv) any data retrieved from tool calls or external databases (rag_main). Some items in the context can be hidden from the user, such as the system prompt or the output of tool calls (zhang2024effective).
@@ 2.1 Private Inference
We define a private inference method for LLMs so that no attacker can reliably infer sensitive information about the input given the output generated by the LLM. Attackers could adaptively query the mechanism, and run membership inference (black_box_system_prompt_extr), or reconstruction attacks (prompt_extraction; inversion_main). In our work, we always assume that all sensitive information is encoded in a subset of tokens in the input. We identify four baseline token-level private inference methods.
1. Scrubbing: Scrubbing is an industry-wide standard often used for removing PII, that relies on modifying the dataset using Named Entity Recognition (NER) to detect sensitive tokens which are redacted or sometimes replaced with private placeholder tokens doc2; doc3. While replacement may not provide perfect privacy, as adjacent tokens can still leak some information (e.g., pronouns leak a person’s gender) staab2024memorizationviolatingprivacyinference, it is widely deployed and accepted as a privatization mechanism.
2. Prompt Engineering. A solution that modifies the inference process is to instruct the model to paraphrase documents without leaking PII mattern2022limits; staab2024large. Compared to NER, this method better preserves the context’s quality, but provides no privacy guarantees. This method cannot be trusted, as (i) previous works showed that it is vulnerable to jailbreak attacks wang2025pig; li2024llm_pbe and (ii) we show that inferential white-box attackers can infer membership at a high success rate without jailbreaking.
3. DP-Decoding: dpdecod proposed DP-decoding, which linearly interpolates the LLM’s output probability distribution (Eq. 1) with a uniform distribution $u$ (i.e., $1/|\mathcal{V}|$ for each token $t$). Then, for a token $y$, the new probability is $\tilde{\Pr}(y\mid x_{<t})=\lambda\,\Pr(y\mid x_{<t})+(1-\lambda)\,u,$ where $\lambda\in[0,1]$ controls the privacy/utility trade-off: larger $\lambda$ allows more of the original LLM distribution to pass, thus improving text quality but reducing privacy (e.g., increasing an attacker’s ability to guess the original input tokens).
4. DP-Prompt: utpala-etal-2023-locally proposed DP-Prompt, which clips the logits ($z$ from Eq. 1) to the range $b_{1},b_{2}]$ and then uses the exponential mechanism to sample the next token $y$. Here, the clipping width $b_{1},b_{2}]$ and the temperature controls the privacy/utility tradeoff.
@@ 2.2 Differential Privacy
Definition 1 (Approximate Differential Privacy (dwork2014algorithmic)).
@@
Here, the parameters $\epsilon>0$ (privacy loss) and $\delta\in[0,1]$ (failure probability) define the privacy guarantee: $\epsilon$ upper bounds the privacy loss, while $\delta$ is the probability that this guarantee does not strictly hold. Stronger privacy corresponds to smaller $\epsilon$ and $\delta$ values. Another notion of DP is called Rényi DP (renyi) that measures privacy loss using the Rényi divergence.
Theorem 1 (Rényi Differential Privacy (RDP) renyi).
@@
$\small D_{\alpha}\!\bigl(\mathsf{M}(D)\,\|\,\mathsf{M}(D^{\prime})\bigr)\leq\epsilon,\qquad D_{\alpha}(P\|Q)=\frac{1}{\alpha-1}\log\mathbb{E}_{x\sim Q}\!\Bigl[\bigl(P(x)/Q(x)\bigr)^{\alpha}\Bigr],$ (3)
@@
Theorem 2 (RDP $\Rightarrow$ DP conversion renyi).
@@
A mechanism is said to satisfy local approximate DP if each individual’s data is randomized on their own device (locally) before transmission, ensuring $(\epsilon,\delta)$-DP with respect to their raw data. Therefore, based on our definition of DPI, DP-Prompt (utpala-etal-2023-locally) is a local pure DPI algorithm (i.e., with $\delta=0$) under a document-level neighborhood. In contrast, DP-Decoding (dpdecod) introduces input-level noise through its output perturbation step, which intuitively provides some privacy for the inference input. However, the original analysis addresses only training data privacy; precise inference time guarantees have not yet been established and remains an open direction for future work.
@@ 3 Threat Model
Adversary’s capabilities and goals. We consider a powerful adaptive gray-box attacker who (i) knows the private inference method used by the defender, (ii) knows the defender’s LLM’s architecture and weights, (iii) observes both the privatized output and (iv) the original document with all private tokens redacted (see 3.a in Figure 2), but does not have access to any hidden activation in the LLM, including logits or final output probabilities from the LLM. The attacker’s objective is to correclty infer the missing sensitive tokens with high probability. We further allow the attacker to access the entire original document, except for the specific privacy group being targeted (e.g., all context except the masked NAME tokens), meaning that all-but-one privacy groups are revealed. This simulates strong attacks that can successfully extract full system prompts (black_box_system_prompt_extr; prompt_extraction). Additionally, we assume that the attacker has a candidate set $C_{j^{*}}$ of possible private tokens, which always includes the true tokens. This interaction is formalized by the following game.
@@ 4 Conceptual Approach
We propose a mechanism with token-level DP guarantees for LLMs during inference inspired by PMixED (flemings2024differentially). PMixED is itself conceptually similar to PATE (Papernot2016SemisupervisedKT) and SUBMIX (ginart2022submix), which aim to achieve DP with respect to records in a training dataset. PATE partitions training data among models and aggregates votes through noisy ensembling, while SUBMIX mixes ensemble predictions with a public pre-trained LM. Instead of adding DP noise with a max-of-Laplacian aggregation mechanism, PMixED uses the inherent stochasticity from sampling LLMs to achieve differential privacy guarantees according to the private prediction framework formalized by pmlr-v75-dwork18a.
In our method, datasets $D$ and $D^{\prime}$ are token sequences in the LLM’s context, where $D^{\prime}$ can be obtained by adding $k$ (private) tokens to $D$. This corresponds to the standard add/remove scheme of DP neighborhood. Our goal is to design a DPI mechanism $\mathcal{A}$ (Defn. 2) to bound the symmetric Rényi divergence ($D_{\alpha}^{\leftrightarrow}$) between $P=\mathcal{A}(D)$ and $Q=\mathcal{A}(D^{\prime})$, such that, $D_{\alpha}^{\leftrightarrow}(P\,\|\,Q)~=~\max\!\Bigl\{\,D_{\alpha}(P\,\|\,Q),~D_{\alpha}(Q\,\|\,P)\Bigr\}$. Where $D_{\alpha}$ is the Rényi divergence (Thm. 1). Our algorithm satisfies $D_{\alpha}^{\leftrightarrow}\!\bigl(\mathcal{A}(D)\,\|\,\mathcal{A}(D^{\prime})\bigr)\leq\alpha\beta$. We use the standard $\alpha=2$, and therefore, $\beta$ is the main controller of privacy-vs-utility and a proxy for $\epsilon$ in our DPI mechanism.
@@ 4.1 DP-FUSION
A complete overview of DP-Fusion is provided in Figure 2. The input is an ordered token sequence separated into privacy groups by an NER oracle. Let there be a token sequence: $D\;=\;(x_{1},x_{2},\dots,x_{N})\;=\;X_{1}\cup X_{2}\cup\dots\cup X_{m}\cup X_{\text{pub}}$ where each token belongs to exactly one privacy group $X_{i}$ ($1\leq i\leq m$) or to the public group $X_{\text{pub}}$, which contains all tokens considered to be non-sensitive. For a Rényi order $\alpha=2$ the user supplies per-group privacy budgets $\beta_{i}$ and the maximum allowed divergence for group $X_{i}$ is therefore $\alpha\beta_{i}$.

### v4

@@ Abstract
Large language models (LLMs) do not preserve privacy at inference-time. The LLM’s outputs can inadvertently reveal information about the model’s context, which presents a privacy challenge when the LLM is augmented via tools or databases containing sensitive information. Existing privacy-preserving methods at inference-time have significant limitations since they (i) lack provable guarantees or (ii) have a poor utility/privacy trade-off. We propose DP-Fusion, a Differentially Private Inference (DPI) mechanism for LLMs that provably bounds the influence a set of tokens in the context can have on the LLM’s output. DP-Fusion works as follows: (1) label a subset of sensitive tokens, (2) infer the LLM without any sensitive tokens to obtain a baseline, (3) infer the LLM with the sensitive tokens, and (4) blend distributions so that the final output remains within a bounded distance of the baseline distribution. While this per-token influence bound also mitigates jailbreak-style prompt injection, we focus on document privatization, where the goal is to paraphrase a document containing sensitive tokens, e.g., personally identifiable information, so that no attacker can reliably infer them from the paraphrased document while preserving high text quality. The privacy/utility trade-off is controlled by $\epsilon$, where $\epsilon=0$ hides sensitive tokens entirely, while higher values trade off privacy for improved text quality. We show that our method creates token-level provably privatized documents with substantially improved theoretical and empirical privacy, achieving $6\times$ lower perplexity than related DPI methods.
Github Repository   |   PyPI Package   |   Deployed Application
@@ 1 Introduction
Figure 1: DP-Fusion differentially private LLM inference.
@@
Large Language Models (LLMs) are trained once and deployed many times. During deployment, LLMs process unseen data they were not trained on, such as user prompts, tool calls or external databases. A privacy challenge emerges when data contains sensitive information such as passwords or Personally Identifiable Information (PII) (Welleck et al., 2024; EU-Regulation, 2016) that the LLM must not reveal to a user. Consider a hospital that wants to deploy LLMs to assist users in matching their symptoms to historical records from a large document dataset.
Many users contributed their doctor’s notes, including health details such as a disease history or a treatment plan linked to PII (Nakka et al., 2024), but expect privacy against re-identification. However, deploying an LLM introduces unique privacy risks since generated tokens could inadvertently and silently leak sensitive data. The challenge is protecting sensitive data while maintaining high service quality.
@@
A straightforward solution would be to carefully label sensitive tokens and scrub them from all documents. Scrubbing is widely applied in practice, but overly aggressive scrubbing has been shown to severely harm utility (Lukas et al., 2023). A better solution could be to fully re-write documents for privacy, e.g., through paraphrasing (Mattern et al., 2022). However, doing inference naively (i) lacks provable guarantees and (ii) our experiments show that attackers can still reliably infer sensitive information when they know which model was used to create the paraphrased text. Private inference solutions fall into two categories: (a) modifying the LLM’s context (e.g., via scrubbing) or (b) modfying the inference process. Dataset-based techniques include randomized methods to replace sensitive tokens in the input (Chen et al., 2022; Tong et al., 2023; Yue et al., 2021). Inference-based techniques modify the model or inference process, e.g., via fine-tuning, prompt engineering (Staab et al., 2024b), adding noise to the output distribution (Majmudar et al., 2022; Utpala et al., 2023). However, existing dataset and inference-based approaches achieve poor privacy/utility trade-offs, either by over-sanitizing the input or by providing weak or no formal guarantees.
@@
We introduce DP-Fusion, a token-level Differentially Private Inference (DPI) method for LLMs that provably bounds the influence of sensitive tokens in the context on generated tokens in its output. Figure˜1 illustrates an overview of our method, which computes the next token to a query on a document containing PII. We first remove the PII from the document, then run the LLM on both the original document (with PII) and the redacted version (without PII). Finally, we mix the probability distributions from both runs, so that the distance between the mixed and original distributions is bounded, sample the next token and return it to the user. Crucially, the attacker’s advantage at inferring the secret is provably bounded even if the query is chosen adversarially (e.g., by selecting a jailbreak attack (Wei et al., 2023)). We empirically demonstrate that our method substantially outperforms all surveyed DPI methods in the utility/privacy trade-off.
@@ 2 Background
An LLM has a context, which typically includes a (i) system prompt, (ii) user queries, (iii) LLM responses and (iv) any data retrieved from tool calls or external databases (Lewis et al., 2020). Some items in the context can be hidden from the user, such as the system prompt or the output of tool calls (Zhang et al., 2024c).
@@ 2.1 Private Inference
We define a private inference method for LLMs so that no attacker can reliably infer sensitive information about the input given the output generated by the LLM. Attackers could adaptively query the mechanism, and run membership inference (Zhang et al., 2024a), or reconstruction attacks (Zhang et al., 2024b; Morris et al., 2023). In our work, we always assume that all sensitive information is encoded in a subset of tokens in the input. We identify four baseline token-level private inference methods.
1. Scrubbing: Scrubbing is an industry-wide standard often used for removing PII, that relies on modifying the dataset using Named Entity Recognition (NER) to detect sensitive tokens which are redacted or sometimes replaced with private placeholder tokens (Mamede et al., 2016; Lison et al., 2021). While replacement may not provide perfect privacy, as adjacent tokens can still leak some information (e.g., pronouns leak a person’s gender) (Staab et al., 2024a), it is widely deployed and accepted as a privatization mechanism.
2. Prompt Engineering. A solution that modifies the inference process is to instruct the model to paraphrase documents without leaking PII (Mattern et al., 2022; Staab et al., 2024b). Compared to NER, this method better preserves the context’s quality, but provides no privacy guarantees. This method cannot be trusted, as (i) previous works showed that it is vulnerable to jailbreak attacks (Wang et al., 2025; Li et al., 2024) and (ii) we show that inferential white-box attackers can infer membership at a high success rate without jailbreaking.
3. DP-Decoding: Majmudar et al. (2022) proposed DP-decoding, which linearly interpolates the LLM’s output probability distribution (Eq. 1) with a uniform distribution $u$ (i.e., $1/|\mathcal{V}|$ for each token $t$). Then, for a token $y$, the new probability is $\tilde{\Pr}(y\mid x_{<t})=\lambda\,\Pr(y\mid x_{<t})+(1-\lambda)\,u,$ where $\lambda\in[0,1]$ controls the privacy/utility trade-off: larger $\lambda$ allows more of the original LLM distribution to pass, thus improving text quality but reducing privacy (e.g., increasing an attacker’s ability to guess the original input tokens).
4. DP-Prompt: Utpala et al. (2023) proposed DP-Prompt, which clips the logits ($z$ from Eq. 1) to the range $b_{1},b_{2}]$ and then uses the exponential mechanism to sample the next token $y$. Here, the clipping width $b_{1},b_{2}]$ and the temperature controls the privacy/utility tradeoff.
@@ 2.2 Differential Privacy
Definition 1 (Approximate Differential Privacy (Dwork et al., 2014)).
@@
Here, the parameters $\epsilon>0$ (privacy loss) and $\delta\in[0,1]$ (failure probability) define the privacy guarantee: $\epsilon$ upper bounds the privacy loss, while $\delta$ is the probability that this guarantee does not strictly hold. Stronger privacy corresponds to smaller $\epsilon$ and $\delta$ values. Another notion of DP is called Rényi DP (Mironov, 2017) that measures privacy loss using the Rényi divergence.
Theorem 1 (Rényi Differential Privacy (RDP) (Mironov, 2017)).
@@
$D_{\alpha}\!\bigl(\mathsf{M}(D)\,\|\,\mathsf{M}(D^{\prime})\bigr)\leq\epsilon,\qquad D_{\alpha}(P\|Q)=\frac{1}{\alpha-1}\log\mathbb{E}_{x\sim Q}\!\Bigl[\bigl(P(x)/Q(x)\bigr)^{\alpha}\Bigr],$ (3)
@@
Theorem 2 (RDP $\Rightarrow$ DP conversion (Mironov, 2017)).
@@
A mechanism is said to satisfy local approximate DP if each individual’s data is randomized on their own device (locally) before transmission, ensuring $(\epsilon,\delta)$-DP with respect to their raw data. Therefore, based on our definition of DPI, DP-Prompt (Utpala et al., 2023) is a local pure DPI algorithm (i.e., with $\delta=0$) under a document-level neighborhood. In contrast, DP-Decoding (Majmudar et al., 2022) introduces input-level noise through its output perturbation step, which intuitively provides some privacy for the inference input. However, the original analysis addresses only training data privacy; precise inference time guarantees have not yet been established and remains an open direction for future work.
@@ 3 Threat Model
Adversary’s capabilities and goals. We consider a powerful adaptive gray-box attacker who (i) knows the private inference method used by the defender, (ii) knows the defender’s LLM’s architecture and weights, (iii) observes both the privatized output and (iv) the original document with all private tokens redacted (see 3.a in Figure 2), but does not have access to any hidden activation in the LLM, including logits or final output probabilities from the LLM. The attacker’s objective is to correclty infer the missing sensitive tokens with high probability. We further allow the attacker to access the entire original document, except for the specific privacy group being targeted (e.g., all context except the masked NAME tokens), meaning that all-but-one privacy groups are revealed. This simulates strong attacks that can successfully extract full system prompts (Zhang et al., 2024a; b). Additionally, we assume that the attacker has a candidate set $C_{j^{*}}$ of possible private tokens, which always includes the true tokens. This interaction is formalized by the following game.
@@ 4 Conceptual Approach
We propose a mechanism with token-level DP guarantees for LLMs during inference inspired by PMixED (Flemings et al., 2024). PMixED is itself conceptually similar to PATE (Papernot et al., 2016) and SUBMIX (Ginart et al., 2022), which target privacy with respect to training-set records—using ensembles over disjoint data partitions and noisy aggregation, PMixED differs by relying on the inherent stochasticity of sampled LLM outputs to provide privacy. SUBMIX adapts PATE to generative modeling, but can exhaust its privacy budget early due to data-dependent accounting, a limitation PMixED overcomes through closed-form Rényi Differential Privacy (RDP) tracking. In contrast to these training-time approaches, our work an approach similar to PMixED into the inference setting under a different threat model: the defender uses an LLM to paraphrase a document while protecting specific sensitive tokens, and the attacker aims to recover those tokens from the paraphrased text. Our method therefore adapts ensemble-style private prediction to operate at inference time, enabling token-level DP guarantees for LLM-generated paraphrases.
In our method, datasets $D$ and $D^{\prime}$ are token sequences in the LLM’s context, where $D^{\prime}$ can be obtained by adding $k$ (private) tokens to $D$. This corresponds to the standard add/remove scheme of DP neighborhood. Our goal is to design a DPI mechanism $\mathcal{A}$ (Defn. 2) to bound the symmetric Rényi divergence ($D_{\alpha}^{\leftrightarrow}$) between $P=\mathcal{A}(D)$ and $Q=\mathcal{A}(D^{\prime})$, such that, $D_{\alpha}^{\leftrightarrow}(P\,\|\,Q)~=~\max\!\Bigl\{\,D_{\alpha}(P\,\|\,Q),~D_{\alpha}(Q\,\|\,P)\Bigr\}$. Where $D_{\alpha}$ is the Rényi divergence (Thm. 1). Our algorithm satisfies $D_{\alpha}^{\leftrightarrow}\!\bigl(\mathcal{A}(D)\,\|\,\mathcal{A}(D^{\prime})\bigr)\leq\alpha\beta$. We use standard $\alpha=2,\delta=0.001$. For a fixed $\alpha$, and $\delta$, number of privacy groups $m$, the resulting $\epsilon$ for the generated tokens is primarily varied by controlling $\beta$ in our DPI mechanism. We observe greater stability when using $\alpha=2$ with regards to the divergence, which can be effectively controlled with small $\lambda$ values, as shown in Appendix A.22. We therefore adopt this setting for all our experiments. This choice is also consistent with prior work in differential privacy, where ($\alpha=2$) is commonly used for simplicity.
@@ 4.1 DP-FUSION
A complete overview of DP-Fusion is provided in Figure 2. The input is an ordered token sequence separated into privacy groups by an NER oracle. Let there be a token sequence: $D\;=\;(x_{1},x_{2},\dots,x_{N})\;=\;X_{1}\cup X_{2}\cup\dots\cup X_{m}\cup X_{\text{pub}}$ where each token belongs to exactly one privacy group $X_{i}$ ($1\leq i\leq m$) or to the public group $X_{\text{pub}}$, which contains all tokens considered to be non-sensitive. To prevent length-based leakage, we pad each redacted span with an equal number of “_” placeholder tokens so that ($X_{\text{pub}}$) and ($X_{\text{pub}}\cup X_{i}$) have identical token lengths. For a Rényi order $\alpha=2$ the user supplies per-group privacy budgets $\beta_{i}$ and the maximum allowed divergence for group $X_{i}$ is therefore $\alpha\beta_{i}$.
@@
DP-Fusion requires $m+1$ forward passes per token (where $m$ is the number of privacy groups) as opposed to $1$ forward pass in the non-private case. However, DP-Fusion is highly parallelizable and the latency is approximately equivalent to that of a single LLM forward pass.
@@ 4.2 Privacy Analysis
Definition 5 (Symmetric Rényi divergence).
For distributions $p,q$ on a common support and Rényi order $\alpha>1$, the symmetric Rényi divergence is $D_{\alpha}^{\leftrightarrow}(p\,\|\,q)=\max\bigl\{D_{\alpha}(p\,\|\,q),\;D_{\alpha}(q\,\|\,p)\bigr\}.$ Under add/remove adjacency (Definition 3), neighboring documents $D\stackrel{{\scriptstyle i}}{{\sim}}D^{\prime}$ require both $D_{\alpha}(M(D)\,\|\,M(D^{\prime}))\leq\alpha\beta_{i}$ and $D_{\alpha}(M(D^{\prime})\,\|\,M(D))\leq\alpha\beta_{i}$. Bounding $D_{\alpha}^{\leftrightarrow}(M(D)\,\|\,M(D^{\prime}))$ enforces these two constraints simultaneously. This divergence is bound in Algorithm 1.